OpenStudy (goformit100):

Find the distance between the lines 3 x+ 4y= 9 and 6x+ 8y= 15.

OpenStudy (anonymous):

3x+ 4y= 9 6x+ 8y= 15 y = -3/4x + 9/4 y = -6/8x + 15/8 y = -3/4x + 9/4 y = -3/4x + 15/8 parallel lines.

OpenStudy (goformit100):

Thank you Sir

OpenStudy (anonymous):

yea, now since we know the lines are parallel, we can just find the distance between where they hit the y axis.

OpenStudy (anonymous):



OpenStudy (anonymous):

distance between the lines = 9/4-15/8

OpenStudy (anonymous):

That's not the distance between the lines. For that we need a perpendicular line to both of them. 4/3 x works. Now we solve this system to see where the intersections are and find the pythagorean distance between them.: \[y=-3x/4+15/8=4x/3 ...x=0.9, y=1.2\] \[y=-3x/4+9/4=4x/3 ...x=1.08, y=1.44\] \[distance=\sqrt{.18^2+.24^2}=.3 \neq9/4-15/8=.375\]